The difference between 2 numbers is 6 Twice the larger added to the smaller is 24 Find the numbers
Answer by jorel1380(3719)

p-n=6
2p+n=24
p=n+6
2(n+6)+n=24
3n+12=24
3n=12
n=4
p=10
The two numbers are 4 and 10..